a toroid is madeof 2 000 turns of wire of radius 2.00 cm formed into a donut shape of inner radius 10.0 cm and the outer radius 14.0 cm. when a 30.0 A current is present to the toroid,determine the magnetic field at a distance of 11.0 cm from the center of the toroid .
1
Expert's answer
2020-05-24T18:05:51-0400
"B=\frac{\mu_0NI}{2\pi r}"

"B=\frac{(4\pi\cdot10^{-7})(2000)(10)}{2\pi (0.11)}=0.0364\ T"